10. The average price of a gallon of gasoline, and the average cost of a barrel of oil, are shown for a five-week period several years ago:

Gasoline ($) Oil ($)
2.089 37.55
2.199 46.20
2.989 77.15
3.009 73.78
3.029 85.78

Calculate the correlation coefficient to four decimal places, and state whether it indicates a strong positive, strong negative, weak positive, weak negative, or no correlation.
Answer #1

.

X Y X⋅Y X⋅X Y⋅Y  
2.089 37.55 78.44195 4.363921 1410.0025
2.199 46.20 101.5938 4.835601 2134.44
2.989 77.15 230.60135 8.934121 5952.1225
3.009 73.78 222.00402 9.054081 5443.4884
3.029 85.78 259.82762 9.174841 7358.2084

\\ Find\ the\ sum\ of\ every\ column \ we\ get\\ {\sum{X}=13.315 ~,~ \sum{Y}=320.46 ~,~ \sum{X \cdot Y}=892.46874 ~,~ \sum{X^2}=36.362565 ~,~ \sum{Y^2}=22298.2618}\\ \\ So,\ correlation\ coefficient \ is\\ \\ \begin{aligned}r~&=~\frac{n\cdot\sum{XY} - \sum{X}\cdot\sum{Y}} {\sqrt{\left[n \sum{X^2}-\left(\sum{X}\right)^2\right] \cdot \left[n \sum{Y^2}-\left(\sum{Y}\right)^2\right]}} \\r~&=~\frac{ 5 \cdot 892.46874 - 13.315 \cdot 320.46 } {\sqrt{\left[ 5 \cdot 36.362565 - 13.315^2 \right] \cdot \left[ 5 \cdot 22298.2618 - 320.46^2 \right] }} \approx 0.9796\end{aligned}\\ \\ Correlation\ coefficient\ is\ positive\ and\ greater\ than\ 0.7\\ So,\ it\ is\ \boldsymbol{{\color{Red} strong \ positive}}\ correlation